OpenStudy (destinyyyy):

Solve the polynomial equation by factoring and then using the zero-product principle.

OpenStudy (destinyyyy):

18x-9=2x^3 -x^2

OpenStudy (destinyyyy):

What?

OpenStudy (destinyyyy):

Whats correct? I posted the question not an answer. And no I cant just type this into a calculator.

OpenStudy (destinyyyy):

36y^3-5=y -180y^2